OpenStudy (anonymous):

What is the factored form of the following expressions? x2 – 10xy + 24y2 A. (x + 6y)(x + 4y) B. (x – 2y)(x + 12y) C. (x + 2y)(x – 12y) D. (x – 6y)(x – 4y)

OpenStudy (anonymous):

I think the answer is b

OpenStudy (anonymous):

Not quite. What are factors of 24 that add up to -10?

OpenStudy (anonymous):

6 and 4

OpenStudy (anonymous):

so the answer is going to be d

OpenStudy (anonymous):

Yes, but just so you know, the factors were -6 an -4 because they add up to -10.

OpenStudy (anonymous):

i know thanks. you going to be around for a while?

OpenStudy (anonymous):

Probably for a little longer.

OpenStudy (anonymous):

can you help me on this next one? 54c3d4 + 9c4d2 A. 9c3d2(d2 + 6c) B. 9c3d2(6d2 + c) C. 9c4d2(d2 + 6) D. 9c4d2(6d2 + 1)

OpenStudy (anonymous):

What is the GCF of 54 and 9?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Ok. What is the GCF of c^3 and c^4?

OpenStudy (anonymous):

c2

OpenStudy (anonymous):

Um no. Think of it this way. c^3 = c * c * c c^4 = c * c * c * c See how there are 3 that can be taken out of both?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How about the GCF of d^4 and d^2?

OpenStudy (anonymous):

d^2

OpenStudy (anonymous):

Alright. You're GCF is 9c^3d^2 \[9c^{3}d^{2}(\frac{54c^{3}d^{4}}{9c^{3}d^{2}} + \frac{9c^{4}d^{2}}{9c^{3}d^{2}})\]

OpenStudy (anonymous):

Can you finish this off?

OpenStudy (anonymous):

the answer is gooing to be b. 9c3d2(6d2 + c)

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)